If z=1.76, Ho Can be rejected at _____.

A) a=.10, but not at a=.05.
B) a=.05 and a=.10
C) a=.05, but not at a=.10
D) a=.01 and a=.05
Answer #1

z=1.76

Assuming that alternative hypothesis is 2 sided

P value=2*P(Z>z)=2*P(Z>1.76)=0.0784

We can reject it at a=0.10 as p value is less than 0.10 but we can't reject it at 0.05

hence answer is

a) a=0.10 but not at a=0.05